The combined age of two people is 75years. In 15years time, one person will be twice the other person's present age. Find their present ages
Answer by scott8148(6628)

x + y = 75

2(x + 15) = y + 15 ___ 2x + 15 = y

substituting ___ x + (2x + 15) = 75 ___ 3x = 60

solve for x, then substitute back to find y